Write the equation of the L satisfying the given geometric condition.

L has y-intercept (0,2) and is perpendicular to the line with equation 2x-3y=6.

Answer by checkley71(8403)

2x-3y=6
-3y=-2x+6
y=-2x/-3+6/-3
y=2x/3-2 this line has a slope (m)=2/3 therefore a perpendicular line would have a slope=-3/2. with a y intercept(b)=2 this line has the equation:
y=-3x/2+2 answer.